OpenStudy (anonymous):

When john and Marsha worked together, they finished a job in ten days. But when she joined him only after he worked for 7 days, they finished the same job in 5 more days. In how many days can each of them finish the job alone?

OpenStudy (dumbcow):

Similar to last problem: The rate they work together is T, the rate that john works is J 10T = 7J + 5T T = (7/5)J 10(7/5)J = 14J takes John 14 days by himself

OpenStudy (anonymous):

how about marsha?

OpenStudy (dumbcow):

Well if M is rate that marsha works T = J + M = 7/5J -> M = (2/5)J 14J = X*M 14J = X*(2/5)J X = 14(5/2) = 35 it takes marsha 35 days